OpenStudy (anonymous):

Linear Algebra: System of Linear Equations. How do I solve for variables in the denominator? Example such as the system of linear equations: (2/x)+ (3/y)+ (4/z)= 1 (9/x)+ (4/y)- (4/z)= 0 (7/x)+ (9/y)- (3/z)= 6

OpenStudy (amistre64):

doesnt matter; same process only instead of =x, you do =1/x

OpenStudy (anonymous):

That is where I am a bit confused. How would I type the augmented matrix?

OpenStudy (amistre64):

2 3 4 1 9 4 4 0 7 9 3 6 id assume

OpenStudy (anonymous):

You assume. Alright, thank you for trying.

OpenStudy (amistre64):

i forgot the negatives :)

OpenStudy (amistre64):

the wolf rrefs it to: http://www.wolframalpha.com/input/?i=rref%7B%7B2%2C+3%2C+4%2C+1%7D%2C%7B9%2C+4%2C+-4%2C+0%7D%2C%7B+7%2C+9%2C+-3%2C+6%7D%7D sooo: x^-1 = -144/257 y^-1 = 263/257 z^-1 = -61/257